OpenStudy (anonymous):

7th grade help please

OpenStudy (anonymous):

Solve. 4 + x/7=2 A) -14 B) 10 C) 12 D) 42

OpenStudy (anonymous):

@Alfa_female

OpenStudy (asnaseer):

I assume you equation is:\[\frac{4+x}{7}=2\]is that right?

OpenStudy (anonymous):

yes